OpenStudy (anonymous):

Help! Use De Moivre's Theorem to compute the following: [3(cos(27)) + isin (27)]^-5

OpenStudy (anonymous):

(1/(3^5))((cos(27*-5)+sin(27*-5))

OpenStudy (anonymous):

Moivre s theorem is easy ; you have to keep the formula in mind !!

OpenStudy (anonymous):

I am new to De Moivre's theorum, okay that equation you said, i continue to solve that and get the answer?

OpenStudy (anonymous):

you have just to replace 27*-5 by -135

OpenStudy (michele_laino):

furthermore, you have to use these identities: \[\begin{gathered} \cos \left( { - 135} \right) = \cos \left( {135} \right) \hfill \\ \sin \left( { - 135} \right) = - \sin \left( {135} \right) \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I continued to solve and got \[-(122 \sqrt(2))/243\]

OpenStudy (anonymous):

@Michele_Laino furthermore he can also use cos(135)=-sin(45) and sin(135)=cos(45)

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Am i correct?

OpenStudy (anonymous):

absolutly noot coorrect !

OpenStudy (anonymous):

Okay I must've did something wrong :/

OpenStudy (anonymous):

you have to find a complex number (a+ib)

OpenStudy (anonymous):

Okay do i plug numbers from the equation into (a + ib)?

OpenStudy (anonymous):

try ! and show me what u get

OpenStudy (anonymous):

Okay Im guessing i is standing for the imaginary number, hmm (27 + i5) ?

OpenStudy (anonymous):

O.O how did u get thiis !

OpenStudy (anonymous):

you have to comput the value of 1/3^5cos(135) and plug it into a

OpenStudy (anonymous):

then compute the value of -1/3^5sin(135) and plug it into b

OpenStudy (anonymous):

Okay for the value should it be a decimal or a fraction?

OpenStudy (anonymous):

a fraction

OpenStudy (anonymous):

take a look ur mail box